OpenStudy (anonymous):

To convert degrees Fahrenheit (F) into degrees Celsius (C) use the formula C=5/9(F-32). Rewrite the equation to convert degrees Celsius to degrees Fahrenheit.

OpenStudy (anonymous):

You want to isolate F on one side of the equation. We could start by multiplying both sides by 9/5 which will eliminate the 5/9 on the right side and get us one step closer to isolating F, so we'll have: \[(9/5)C=F-32\]Now, to get F by itself we just need to add 32 to both sides: \[(9/5)C + 32=F\]